[Brahms-l] Shift report 20040123 16:00-24:00

From: Apache <apache@bnl.gov>
Date: Fri Jan 23 2004 - 23:51:46 EST
------------------------------------------------------------
Date: Fri Jan 23 23:51:46 2004, Shift: 16:00-24:00
Supervisor: Ionut ARSENE, with: Erik JOHNSON
Log book#15 pages 41 - 41

Shift Summary:
------------------------------------------------------------
1. Beam for almost the entire shift.

2. Pedestals:
 9834 - Non-TPC pedestals
 9835 - TPC pedestals
   Data:
 settings FS@10deg with D1@B430 && MRS@40deg with D5@B1050
 9829,9830 - 380k trig6 & 320k trig3 (this setting is DONE)
 settings FS@10deg with D1@B608 && MRS@40deg with D5@B1050
 9831, 9832, 9833 - 430k trig6 & 430k trig3
 9836(still running) - ~250k trig6 & ~220k trig3

3. No problems detected. 
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Fri Jan 23 23:52:16 2004

This archive was generated by hypermail 2.1.8 : Fri Jan 23 2004 - 23:52:34 EST